OpenStudy (anonymous):

if a car can fo from 0.0 to 60.0mi/hr in 8 seconds, what would be its final speed after 5.o seconds if its starting speed were 50.o mi/hr?

OpenStudy (anonymous):

Acceleration = delta velocity/delta time = 60-0/8 seconds = 7.5 mi/hr/s 50 + 7.5(5) = 87.5 mi/hr